# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 1. Introduction
Let $n$ be a positive integer. We denote by $\mathbb{F}_{q}$ the finite field with $q=2^{n}$ elements, by $\mathbb{F}_{q}^{*}$ the multiplicative cyclic group of non-zero elements of $\mathbb{F}_{q}$ and by $\mathbb{F}_{q}[X]$ the ring of polynomials in one variable $X$ with coefficients in $\mathbb{F}_{q}$. For a function $F:\mathbb{F}_{q}\to\mathbb{F}_{q}$, and for any $a\in\mathbb{F}_{q}$, the derivative of $F$ in the direction $a$ is defined as $D_{F}(X,a):=F(X+a)+F(X)$ for all $X\in\mathbb{F}_{q}$. For any $a,b\in\mathbb{F}_{q}$, the Difference Distribution Table (DDT) entry $\operatorname{DDT}_{F}(a,b)$ at point $(a,b)$ is the number of solutions $X\in\mathbb{F}_{q}$ of the equation $D_{F}(X,a)=b$. Further, the differential uniformity of $F$, denoted by $\Delta_{F}$, is given by $\Delta_{F}:=\max\left\{\operatorname{DDT}_{F}(a,b)\bigm{|}a\in\mathbb{F}_{q}^{*},b\in\mathbb{F}_{q}\right\}.$ We call a function $F$ an almost perfect nonlinear (APN) function if $\Delta_{F}=2$. Notice that $\Delta_{F}$ is always even, because if $X$ is a solution of $D_{F}(X,a)=b$ then $X+a$ is also a solution of $D_{F}(X,a)=b$. Functions with low differential uniformity are more resistant to differential attacks, making them desirable for cryptographic applications.
@@
Recently, Eddahmani and Mesnager [12] studied various properties of the EBCT, LBCT and UBCT. They also determined these entries for the inverse function. In a separate work, Man et al. [16] also gave the DBCT, LBCT and UBCT entries for the inverse function. In this work, we discuss the DBCT entries of the Gold function, and we provide a connection between these new EBCT, LBCT and UBCT concepts and the DDT entries for a differentially $\delta$-uniform function. As a consequence, it turns out that our result covers a particular case of a result of Eddahmani et al. [12] and Man et al. [16] for the inverse function over $\mathbb{F}_{2^{n}}$. Moreover, we explicitly compute EBCT, LBCT and UBCT entries of three classes of differentially-4 uniform power permutations. We shall now give the structure of the paper. We first recall some definitions and results in Section 2. In Section 3, we provide some general results on the EBCT. We also give a generalization for the definitions of the EBCT, LBCT and UBCT for any function over $\mathbb{F}_{2^{n}}$. In Section 4, we discuss the invariance of the EBCT, LBCT and UBCT under the CCZ-equivalence. In Section 5, we find a connection between the EBCT, LBCT and UBCT entries of a differentially $\delta-$uniform function and also express them in terms of the DDT entries. Section 6 and Section 7 deal with the EBCT, LBCT and UBCT entries of APN functions (leading to a characterization of APN functions) and 4-differential uniform functions, as well as some other well-known functions. We also computed the DBCT entries for the Gold function in Section 8. Finally, we conclude the paper in Section 9.
@@ 2. Preliminaries
We will first recall several definitions and lemmas used in the subsequent sections.
@@
Recently, Eddahmani and Mesnager [12] and, Man et al. [16] redefined the notions of LBCT and UBCT for a permutation $F$ without involving its compositional inverse, as stated in the following lemmas.
@@
$\operatorname{LBCT}_{F}(a,b,c)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}X+Y=b\\ F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\end{cases}\right\}\right\rvert,$
@@
$\operatorname{UBCT}_{F}(a,b,c)=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\\ F(X)+F(X+a)=b\end{cases}\right\}\right\rvert.$
@@
(Feistel Boomerang Connectivity Table) [3] Let $F:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ and $a,b\in\mathbb{F}_{2^{n}}$. The Feistel Boomerang Connectivity Table ($\operatorname{FBCT}$) of $F$ is given by a $2^{n}\times 2^{n}$ table, of entry at $(a,b)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ defined by
@@
Moreover, in [12] it was shown that for $a,b,c,d\in\mathbb{F}_{2^{n}}$ with $abcd=0$, the values of $\operatorname{EBCT}_{F}(a,b,c,d)$ are trivial to compute as are the values of $\operatorname{LBCT}_{F}(a,b,c)$ and $\operatorname{UBCT}_{F}(a,b,c)$, for $abc=0$. We recall these results in the following three lemmas.
@@
$\operatorname{EBCT}_{F}(a,b,c,d)$ =|$=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}2^{n}&\mbox{if}~{}a=b=c=d=0\\ \operatorname{DDT}_{F}(c,d)&\mbox{if}~{}a=b=0,c\neq 0,d\neq 0\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}a\neq 0,b\neq 0,c=d=0\\ 0&\mbox{otherwise}\end{cases}\right\}\right\rvert,$
$\operatorname{LBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}2^{n}&\mbox{if}~{}b=c=0\\ 0&\mbox{if}~{}b\neq 0,c=0\\ \operatorname{DDT}_{F}(b,c)&\mbox{if}~{}a=0\\ 0&\mbox{if}~{}b=0,c\neq 0\end{cases}\right\}\right\rvert,$
$\operatorname{UBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}2^{n}&\mbox{if}~{}a=b=0\\ 0&\mbox{if}~{}a=0,b\neq 0,c\neq 0\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}c=0\\ 0&\mbox{if}~{}a\neq 0,b=0,c\neq 0\end{cases}\right\}\right\rvert.$
Also, Wagner [18] gives a relation between UBCT and DDT as $\operatorname{UBCT}_{F}(a,b,b)=\operatorname{DDT}_{F}(a,b)$ and between UBCT and BCT as $\text{BCT}_{F}(a,c)=\sum_{b\in\mathbb{F}_{2^{n}}}\operatorname{UBCT}_{F}(a,b,c)$. Additionally, it is easy to see that $\operatorname{DDT}_{F}(a,c)=\operatorname{LBCT}_{F}(a,a,c)$. Next, we recall the notion of DBCT as follows.
@@ 3. General results on the EBCT
As a consequence, we can derive the connection between the EBCT of a function and the EBCT of its compositional inverse.
@@
$\operatorname{EBCT}_{F}(a,b,c,d)$ =|$=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X)+F(X+a)=b\\ F(X)+F(X+c)=d\\ F(X+a+c)+F(X+a)=d\end{cases}\right\}\right\rvert,$
$\operatorname{LBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}X+Y=b\\ F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\end{cases}\right\}\right\rvert,$
$\operatorname{UBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{(X,Y)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\\ F(X)+F(X+a)=b\end{cases}\right\}\right\rvert.$
@@
Note that if $F$ is a permutation, these definitions are equivalent to the original definitions of these concepts. Therefore, we will refer to this definition for all functions.
@@ 4. Invariance under the CCZ, extended affine and affine-equivalence
We recall that two functions $F,G:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{m}}$ are CCZ-equivalent [6] if there exists an affine permutation ${\mathcal{A}}$ on $\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{m}}$ such that
$\left\{\begin{pmatrix}x\\ G(x)\end{pmatrix},x\in\mathbb{F}_{2^{n}}\right\}=\left\{{\mathcal{A}}\begin{pmatrix}x\\ F(x)\end{pmatrix},x\in\mathbb{F}_{2^{n}}.\right\}.$
As customary, we use the natural identification of the elements in $\mathbb{F}_{2^{n}}$ with the elements in $\mathbb{F}_{2}^{n}$, and, by abuse, we denote by $x$ both an element in $\mathbb{F}_{2^{n}}$ and the corresponding element in $\mathbb{F}_{2}^{n}$. We also decompose the affine permutation ${\mathcal{A}}$ as an affine block-matrix, for an input vector $u\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{m}}$: ${\mathcal{A}}{u}=\begin{pmatrix}{\mathcal{A}}_{11}&{\mathcal{A}}_{12}\\ {\mathcal{A}}_{21}&{\mathcal{A}}_{22}\end{pmatrix}{u}+\begin{pmatrix}C\\ D\end{pmatrix}$, where ${\mathcal{A}}_{11}:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}},{\mathcal{A}}_{12}:\mathbb{F}_{2^{m}}\to\mathbb{F}_{2^{n}}$, ${\mathcal{A}}_{21}:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{m}},{\mathcal{A}}_{22}:\mathbb{F}_{2^{m}}\to\mathbb{F}_{2^{m}}$ and $\begin{pmatrix}C\\ D\end{pmatrix}$ is a column vector in $\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{m}}$. For us, below, $m=n$.
When ${\mathcal{A}}_{12}=0$, we say that $F$ and $G$ are Extended Affine (EA)-equivalent. This can be also written as $G=F_{2}\circ F\circ F_{1}+F_{0}$, for some $F_{i}(X)=L_{i}(X)+B_{i}$ affine functions, where $L_{i}$ are linearized polynomials, $B_{i}\in\mathbb{F}_{2^{n}}$, and $F_{1}$ and $F_{2}$ are permutation polynomials.
When ${\mathcal{A}}_{12}={\mathcal{A}}_{21}=0$, we say that $F$ and $G$ are Affine-equivalent. This can be also written as $G=F_{2}\circ F\circ F_{1}$, for some $F_{i}(X)=L_{i}(X)+B_{i}$ affine functions, where $L_{i}$ are linearized polynomials, $B_{i}\in\mathbb{F}_{2^{n}}$, and $F_{1}$ and $F_{2}$ are permutation polynomials.
@@
Taking $z=y+\gamma$ and $\delta={\mathcal{A}}_{22}d+{\mathcal{A}}_{21}c$, we obtain:
$G(y+\gamma)+G(y)=\delta.$
We can also see that
@@
Therefore,
(4.1) $F(X)+F(X+c)=d\Rightarrow G(y+\gamma)+G(y)=\delta,$
where $\gamma={\mathcal{A}}_{12}d+{\mathcal{A}}_{11}c\mbox{ and }\delta={\mathcal{A}}_{22}d+{\mathcal{A}}_{21}c.$
Similarly, taking $\begin{pmatrix}u\\ G(u)\end{pmatrix}={\mathcal{A}}\begin{pmatrix}X+a\\ F(X+a)\end{pmatrix}$, we can prove that
(4.2) $F(X)+F(X+a)=b\Rightarrow G(y+\alpha)+G(y)=\beta,$
where $\alpha={\mathcal{A}}_{12}b+{\mathcal{A}}_{11}a\mbox{ and }\beta={\mathcal{A}}_{22}b+{\mathcal{A}}_{21}a$.
@@
Therefore, we get the desired equation
$G(y+\alpha)+G(y+\alpha+\gamma)=\delta,$
with the same values of $\alpha,\gamma,$ and $\delta$. This concludes the proof. ∎
@@
The LBCT spectrum is preserved under EA-equivalence, but not under CCZ-equivalence. If $F$ and $G$ are EA-equivalent, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{LBCT}_{G}(\alpha_{L},\beta_{L},\gamma_{L})$, where $\alpha_{L}={\mathcal{A}}_{11}a$,$\beta_{L}={\mathcal{A}}_{11}b\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b$.
NB: If $F$ is APN, then CCZ and EA equivalences are equivalent, and therefore, trivially, in this case the LBCTF spectrum is preserved under the CCZ-equivalence.
@@
(4.3) $F(X)+F(X+b)=c\Rightarrow G(y+\beta_{L})+G(y)=\gamma_{L},$
@@
we obtain
(4.4) $F(X+a)+F(X+a+b)=c\Rightarrow G(y+\alpha_{L})+G(y+\alpha_{L}+\beta_{L})=\gamma_{L},$
where $\alpha_{L}={\mathcal{A}}_{12}(F(X)+F(X+a))+{\mathcal{A}}_{11}a\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b,\,\beta_{L}={\mathcal{A}}_{12}c+{\mathcal{A}}_{11}b$.
We see that, unless $F$ is APN, $\alpha_{L}$ is not a constant. Therefore, for general $F$, we have that the spectrum of $\operatorname{LBCT}_{F}(a,b,c)$ is invariant under ${\mathcal{A}}$ if and only if ${\mathcal{A}}_{12}=0$, which gives EA-equivalence. If $F$ and $G$ are EA-equivalent, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{LBCT}_{G}(\alpha_{L},\beta_{L},\gamma_{L})$, where $\alpha_{L}={\mathcal{A}}_{11}a$,$\beta_{L}={\mathcal{A}}_{11}b\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b$. ∎
@@
The UBCT spectrum is preserved under affine equivalence, but it is not preserved under CCZ or EA-equivalence. If $F$ and $G$ are affine equivalent, then $\operatorname{UBCT}_{F}(a,b,c)=\operatorname{UBCT}_{G}(\alpha_{U},\beta_{U},\gamma_{U})$, where $\alpha_{U}={\mathcal{A}}_{11}a,\,\beta_{U}={\mathcal{A}}_{22}b\mbox{ and }\gamma_{U}={\mathcal{A}}_{22}c$.
@@
(4.5) $F(X)+F(X+a)=b\Rightarrow G(y+\alpha_{U})+G(y)=\beta_{U},$

### v3

@@ 1. Introduction
Let $n$ be a positive integer. We denote by $\mathbb{F}_{2^{n}}$ the finite field with $2^{n}$ elements, by $\mathbb{F}_{2^{n}}^{*}$ the multiplicative cyclic group of non-zero elements of $\mathbb{F}_{2^{n}}$ and by $\mathbb{F}_{2^{n}}[X]$ the ring of polynomials in one variable $X$ with coefficients in $\mathbb{F}_{2^{n}}$. For a function $F:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$, and for any $a\in\mathbb{F}_{2^{n}}$, the derivative of $F$ in the direction $a$ is defined as $D_{F}(X,a):=F(X+a)+F(X)$ for all $X\in\mathbb{F}_{2^{n}}$. For any $a,b\in\mathbb{F}_{2^{n}}$, the Difference Distribution Table (DDT) entry $\operatorname{DDT}_{F}(a,b)$ at point $(a,b)$ is the number of solutions $X\in\mathbb{F}_{2^{n}}$ of the equation $D_{F}(X,a)=b$. Further, the differential uniformity of $F$, denoted by $\Delta_{F}$, is given by $\Delta_{F}:=\max\left\{\operatorname{DDT}_{F}(a,b)\bigm{|}a\in\mathbb{F}_{2^{n}}^{*},b\in\mathbb{F}_{2^{n}}\right\}.$ We call a function $F$ an almost perfect nonlinear (APN) function if $\Delta_{F}=2$. Notice that $\Delta_{F}$ is always even, because if $X$ is a solution of $D_{F}(X,a)=b$ then $X+a$ is also a solution of $D_{F}(X,a)=b$. Functions with low differential uniformity are more resistant to differential attacks, making them desirable for cryptographic applications.
@@
Recently, Eddahmani and Mesnager [13] studied various properties of the EBCT, LBCT and UBCT. They also determined these entries for the inverse function. In a separate work, Man et al. [17] also gave the DBCT, LBCT and UBCT entries for the inverse function. In this work, we discuss the DBCT entries of the Gold function, and we provide a connection between these new EBCT, LBCT and UBCT concepts and the DDT entries for a differentially $\delta$-uniform function. As a consequence, it turns out that our result covers a particular case of a result of Eddahmani et al. [13] and Man et al. [17] for the inverse function over $\mathbb{F}_{2^{n}}$. Moreover, we explicitly compute EBCT, LBCT and UBCT entries of three classes of differentially-4 uniform power permutations.
We shall now give the structure of the paper. We first recall some definitions and results in Section 2. In Section 3, we provide some general results on the EBCT. We also give a generalization for the definitions of the EBCT, LBCT and UBCT for any function over $\mathbb{F}_{2^{n}}$. In Section 4, we discuss the invariance of the EBCT, LBCT and UBCT under the CCZ-equivalence. In Section 5, we find a connection between the EBCT, LBCT and UBCT entries of a differentially $\delta-$uniform function and also express them in terms of the DDT entries. Section 6 and Section 7 deal with the EBCT, LBCT and UBCT entries of APN functions (leading to a characterization of APN functions) and 4-differential uniform functions, as well as some other well-known functions. We also computed the DBCT entries for the Gold function in Section 8. Finally, we conclude the paper in Section 9.
@@ 2. Preliminaries
We will first recall several definitions and lemmas used in the subsequent sections. Throughout the paper, we shall use ${\rm Tr}_{m}^{n}$ to denote the (relative) trace function from $\mathbb{F}_{2^{n}}\rightarrow\mathbb{F}_{2^{m}}$, i.e., ${\rm Tr}_{m}^{n}(X)=\sum_{i=0}^{\frac{n-m}{m}}X^{2^{mi}}$, where $m$ and $n$ are positive integers and $m\mid n$. For $m=1$, we use ${\rm Tr}$ to denote the absolute trace. As customary, for a set $S$, we let $a+S=\{a+s\,|\,s\in S\}$.
@@
Recently, Eddahmani and Mesnager [13] and, Man et al. [17] redefined the notions of LBCT and UBCT for a permutation $F$ without involving its compositional inverse ([17] also mentions the non-necessity of the permutation property for the UBCT), as stated in the following lemmas.
@@
$\operatorname{LBCT}_{F}(a,b,c)=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists~{}Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}X+Y=b\\ F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\end{cases}\right\}\right\rvert,$
@@
$\operatorname{UBCT}_{F}(a,b,c)=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists~{}Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\\ F(X)+F(X+a)=b\end{cases}\right\}\right\rvert.$
@@
(Feistel Boomerang Connectivity Table) [3] Let $F:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ and $a,b\in\mathbb{F}_{2^{n}}$. The Feistel Boomerang Connectivity Table (FBCT) of $F$ is given by a $2^{n}\times 2^{n}$ table, of entry at $(a,b)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ defined by
@@
Moreover, in [13] it was shown that for $a,b,c,d\in\mathbb{F}_{2^{n}}$ with $abcd=0$, the values of $\operatorname{EBCT}_{F}(a,b,c,d)$ are simple to compute if $F$ is a permutation, as are the values of $\operatorname{LBCT}_{F}(a,b,c)$ and $\operatorname{UBCT}_{F}(a,b,c)$, for $abc=0$. We recall these results in the following three lemmas.
@@
$\operatorname{EBCT}_{F}(a,b,c,d)$ =$=\begin{cases}2^{n}&\mbox{if}~{}a=b=c=d=0,\\ \operatorname{DDT}_{F}(c,d)&\mbox{if}~{}a=b=0,c\neq 0,d\neq 0,\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}a\neq 0,b\neq 0,c=d=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{LBCT}_{F}(a,b,c)$ =$=\begin{cases}2^{n}&\mbox{if}~{}b=c=0,\\ 0&\mbox{if}~{}b\neq 0,c=0,\\ \operatorname{DDT}_{F}(b,c)&\mbox{if}~{}a=0,\\ 0&\mbox{if}~{}b=0,c\neq 0,\end{cases}$
$\operatorname{UBCT}_{F}(a,b,c)$ =$=\begin{cases}2^{n}&\mbox{if}~{}a=b=0,\\ 0&\mbox{if}~{}a=0,b\neq 0,c\neq 0,\\ \operatorname{DDT}_{F}(a,b)&\mbox{if}~{}c=0,\\ 0&\mbox{if}~{}a\neq 0,b=0,c\neq 0.\end{cases}$
Remark 2.9.
If $F$ is not a permutation, some of the results of Lemma 2.8 are not true. We give the generalization of Lemma 2.8 in Section 5.
Also, Wagner [19] shows that $\operatorname{UBCT}_{F}(a,b,b)=\operatorname{DDT}_{F}(a,b)$, as well as, $\text{BCT}_{F}(a,c)=\sum_{b\in\mathbb{F}_{2^{n}}}\operatorname{UBCT}_{F}(a,b,c)$. Additionally, it is easy to see that $\operatorname{DDT}_{F}(a,c)=\operatorname{LBCT}_{F}(a,a,c)$. Next, we recall the notion of DBCT as follows.
@@ 3. General results on the EBCT
As a consequence, we can derive the connection between the EBCT of a permutation and the EBCT of its compositional inverse.
@@
(3.1) $\operatorname{EBCT}_{F}(a,b,c,d)$ =|$=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\begin{cases}F(X)+F(X+a)=b\\ F(X)+F(X+c)=d\\ F(X+a+c)+F(X+a)=d\end{cases}\right\}\right\rvert,$
(3.2) $\operatorname{LBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}X+Y=b\\ F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\end{cases}\right\}\right\rvert,$
(3.3) $\operatorname{UBCT}_{F}(a,b,c)$ =|$=\left\lvert\left\{X\in\mathbb{F}_{2^{n}}\bigg{|}\exists Y\in\mathbb{F}_{2^{n}}\mbox{ with }\begin{cases}F(X+a)+F(Y+a)=c\\ F(X)+F(Y)=c\\ F(X)+F(X+a)=b\end{cases}\right\}\right\rvert.$
@@
Note that if $F$ is a permutation, these definitions are equivalent to the original definitions of these concepts. We preferred to define these concepts this way to avoid overcounting, but we point out that even if we count the number of pairs $(X,Y)$ in each set, all results hold, except for the UBCT in Theorem 5.1, where its expression becomes too cumbersome.
@@ 4. Invariance under the CCZ, extended affine and affine-equivalence
We recall that two functions $F,G:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ are CCZ-equivalent [7] if there exists an affine permutation ${\mathcal{A}}$ on $\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ such that
$\left\{\begin{pmatrix}x\\ G(x)\end{pmatrix}\,|\,x\in\mathbb{F}_{2^{n}}\right\}=\left\{{\mathcal{A}}\begin{pmatrix}x\\ F(x)\end{pmatrix}\,|\,x\in\mathbb{F}_{2^{n}}\right\}.$
As customary, we use the natural identification of the elements in $\mathbb{F}_{2^{n}}$ with the elements in $\mathbb{F}_{2}^{n}$, and, by abuse, we denote by $x$ both an element in $\mathbb{F}_{2^{n}}$ and the corresponding element in $\mathbb{F}_{2}^{n}$. We also decompose the affine permutation ${\mathcal{A}}$ as an affine block-matrix, for an input vector $u\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$: ${\mathcal{A}}{u}=\begin{pmatrix}{\mathcal{A}}_{11}&{\mathcal{A}}_{12}\\ {\mathcal{A}}_{21}&{\mathcal{A}}_{22}\end{pmatrix}{u}+\begin{pmatrix}C\\ D\end{pmatrix}$, where ${\mathcal{A}}_{ij}:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}},i,j\in\{1,2\}$ and $\begin{pmatrix}C\\ D\end{pmatrix}$ is a column vector in $\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$.
When ${\mathcal{A}}_{12}=0$, we say that $F$ and $G$ are extended affine (EA)-equivalent. This can be also written as $G=F_{2}\circ F\circ F_{1}+F_{0}$, for some $F_{i}(X)=L_{i}(X)+B_{i}$ affine functions, where $L_{i}$ are linearized polynomials, $B_{i}\in\mathbb{F}_{2^{n}}$ such that $F_{1}$ and $F_{2}$ are permutation polynomials.
When ${\mathcal{A}}_{12}={\mathcal{A}}_{21}=0$, we say that $F$ and $G$ are affine-equivalent. This can be also written as $G=F_{2}\circ F\circ F_{1}$, for some $F_{i}(X)=L_{i}(X)+B_{i}$ affine functions, where $L_{i}$ are linearized polynomials, $B_{i}\in\mathbb{F}_{2^{n}}$ such that $F_{1}$ and $F_{2}$ are permutation polynomials.
Here, we discuss the behavior of the EBCT, LBCT and UBCT of a function $F$ under CCZ, extended affine and affine-equivalence. It is worth noting that Eddahmani and Mesnager [13] proved that the EBCT, LBCT and UBCT of a permutation remain invariant under affine-equivalence.
@@
Taking $z=y+\gamma$ and $\delta={\mathcal{A}}_{22}d+{\mathcal{A}}_{21}c$, we obtain: $G(y+\gamma)+G(y)=\delta.$ We can also see that
@@
Therefore, $F(X)+F(X+c)=d\Rightarrow G(y+\gamma)+G(y)=\delta,$ where $\gamma={\mathcal{A}}_{12}d+{\mathcal{A}}_{11}c\mbox{ and }\delta={\mathcal{A}}_{22}d+{\mathcal{A}}_{21}c.$
Similarly, taking $\begin{pmatrix}u\\ G(u)\end{pmatrix}={\mathcal{A}}\begin{pmatrix}X+a\\ F(X+a)\end{pmatrix}$, we can prove that $F(X)+F(X+a)=b\Rightarrow G(y+\alpha)+G(y)=\beta,$ where $\alpha={\mathcal{A}}_{12}b+{\mathcal{A}}_{11}a\mbox{ and }\beta={\mathcal{A}}_{22}b+{\mathcal{A}}_{21}a$.
@@
Therefore, we get the desired equation $G(y+\alpha)+G(y+\alpha+\gamma)=\delta,$ with the same values of $\alpha,\gamma,$ and $\delta$.
Finally, we note that
$\left(\begin{array}[]{c}\alpha\\ \beta\\ \gamma\\ \delta\end{array}\right)=\left(\begin{array}[]{cccc}{\mathcal{A}}_{11}&{\mathcal{A}}_{12}&0&0\\ {\mathcal{A}}_{21}&{\mathcal{A}}_{22}&0&0\\ 0&0&{\mathcal{A}}_{11}&{\mathcal{A}}_{12}\\ 0&0&{\mathcal{A}}_{21}&{\mathcal{A}}_{22}\end{array}\right)\left(\begin{array}[]{c}a\\ b\\ c\\ d\end{array}\right)=\mathcal{B}\left(\begin{array}[]{c}a\\ b\\ c\\ d\end{array}\right)$
and, since ${\mathcal{A}}$ is invertible, the matrix $\mathcal{B}$ is also invertible. This concludes the proof. ∎
@@
The LBCT spectrum is preserved under EA-equivalence. If $F$ and $G$ are EA-equivalent, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{LBCT}_{G}(\alpha_{L},\beta_{L},\gamma_{L})$, where $\alpha_{L}={\mathcal{A}}_{11}a$,$\beta_{L}={\mathcal{A}}_{11}b\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b$.
@@
$F(X)+F(X+b)=c\Rightarrow G(y+\beta_{L})+G(y)=\gamma_{L},$
@@
we obtain $F(X+a)+F(X+a+b)=c\Rightarrow G(y+\alpha_{L})+G(y+\alpha_{L}+\beta_{L})=\gamma_{L},$ where $\alpha_{L}={\mathcal{A}}_{12}(F(X)+F(X+a))+{\mathcal{A}}_{11}a\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b,\,\beta_{L}={\mathcal{A}}_{12}c+{\mathcal{A}}_{11}b$.
We see that, in general, $\alpha_{L}$ is not a constant. However, if ${\mathcal{A}}_{12}=0$ (i.e. EA-equivalence), we have that $\alpha_{L}$ is a constant. Therefore, if $F$ and $G$ are EA-equivalent, then $\operatorname{LBCT}_{F}(a,b,c)=\operatorname{LBCT}_{G}(\alpha_{L},\beta_{L},\gamma_{L})$, where $\alpha_{L}={\mathcal{A}}_{11}a$,$\beta_{L}={\mathcal{A}}_{11}b\mbox{ and }\gamma_{L}={\mathcal{A}}_{22}c+{\mathcal{A}}_{21}b$.
Finally,
$\left(\begin{array}[]{c}\alpha_{L}\\ \beta_{L}\\ \gamma_{L}\end{array}\right)=\left(\begin{array}[]{ccc}{\mathcal{A}}_{11}&0&0\\ 0&{\mathcal{A}}_{11}&0\\ 0&{\mathcal{A}}_{21}&{\mathcal{A}}_{22}\end{array}\right)\left(\begin{array}[]{c}a\\ b\\ c\end{array}\right)={\mathcal{C}}\left(\begin{array}[]{c}a\\ b\\ c\end{array}\right),$
and, since ${\mathcal{A}}$ is invertible, the matrix ${\mathcal{C}}$ is also invertible. This concludes the proof. ∎
@@
The UBCT spectrum is preserved under affine equivalence. If $F$ and $G$ are affine equivalent, then $\operatorname{UBCT}_{F}(a,b,c)=\operatorname{UBCT}_{G}(\alpha_{U},\beta_{U},\gamma_{U})$, where $\alpha_{U}={\mathcal{A}}_{11}a,\,\beta_{U}={\mathcal{A}}_{22}b\mbox{ and }\gamma_{U}={\mathcal{A}}_{22}c$.
@@
$F(X)+F(X+a)=b\Rightarrow G(y+\alpha_{U})+G(y)=\beta_{U},$